A brick mason requires 12 hours to lay a garden wall by himself. After the mason and his helper worked for 3 hours, the mason left to finish a different job. The helper took 12 more hours to finish the job. How long would it have taken for the helper working alone to do the whole job?
Answer by richwmiller(17219)

One worker works alone for 12/x
Two workers work together for 3/x and 3/12 since he can do the job alone in 12
One worker worked 3 hours with his partner and 12 hours alone for a totel of 15 hours
12/x+3/x+3/12=1
15/x+3/12=1
15/x+=1-3/12
15/x+=0.75
15=0.75*x
15/0.75=x
x=20 hours if he had done the job alone
check
12/20+3/20+3/12=1
0.6+0.15+0.25=1
1=1
ok